OpenStudy (anonymous):

The seventh term of an arithmetic sequence is 21, and the tenth term is 126. The first term is _____.

OpenStudy (anonymous):

-189??

myininaya (myininaya):

yes that is right! :)

myininaya (myininaya):

Like this is how I would look at it: we have these numbers: \[a_1,a_2,a_3,a_4,a_5,a_6,a_7,a_8,a_9,a_{10}\] And a_7=21 a_10=126 so we have \[a_1,a_2,a_3,a_4,a_5,a_6,21,a_8,a_9,126\] \[21-6x,21-5x,21-4x,21-3x,21-2x,21-x,21,21+x,21+2x,21+3x=126\] So we can solve that last equation for x

myininaya (myininaya):

21+3x=126 x=105/3=35 And then take that number and plug into the first term